A small business company is considering updating the current production line. There are two plans. For plan A, the fixed cost will be $40,000 and the variable cost will be $27 per unit after the update. For plan B, the fixed costs will be $51,000 and the variable cost will be $26 per unit after the update. Please answer the following question:

Suppose the selling price is $32. Also, the company aims to achieve a profit of $9,000 after the update. What selling volume will be required to achieve the profit for each plan? Which plan has a lower volume?
Answer #1

Plan A

Let the volume in order to obtain the profit be X, then

Profit = Revenue - Expense

Profit = Selling price * Volume - (Fixed cost + Variable cost)

9000 = 32*X - (40000 + 27*X)

5*X = 49000

X = 9800 Units

Plan B

Let the volume in order to obtain the profit be Y, then

Profit = Selling price*Volume - (Fixed cost + Variable cost)

9000 = 32*Y - (51000+26*Y)

6*Y = 60000

Y = 10000 Units

Plan A has a lower volume based on the data.